OpenStudy (anonymous):

rewrite in terms of sin x cos x. Sin 4x

OpenStudy (anonymous):

sin (4x) = sin (2x + 2x) = sin (2x) cos(2x) + sin (2x)cos(2x) = 2 sin(2x)cos(2x) I used the formula for sin (A+B) = sin A cos B + sin B cos A so A = 2x and B = 2x in this problem

OpenStudy (anonymous):

@Easyaspi314 is right but you need to again expand sin 2x and cos 2x using formulae sin(x+x) and cos(x+x)

OpenStudy (anonymous):

OpenHware..I dont believe that was the intent of the question...as cos(2x) will never be just in six or cos x...best scenario is cos^2x or sin^2x

OpenStudy (anonymous):

yeah (cos x) ^2 and (sin x)^2 are still in terms of cos x and sin x.............and I believe the question was to breakdown the given expression in terms of cos x and sin x

OpenStudy (anonymous):

The problem didnt specify. So I think it is fine.

OpenStudy (anonymous):

We can't split hairs or second-guess without specifically stating sin x or/and cos x ONLY.

OpenStudy (anonymous):

and cos(2x) is technically also in terms of cos x, just like cos^2(x) is in terms of cos x.